OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 2(x + 6) - 10 = 12 + 4(x - 1) In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

hero (hero):

If it were me, I'd probably make a list of statements and reasons

hero (hero):

for each step

OpenStudy (anonymous):

you want to get all the numbers to one side of the equals sign and all the letters to the right, so first you want to get rid of the brackets. so the equation would be 2x+12-10=12+4x-4

OpenStudy (anonymous):

sorry all the letters to the other side, doesnt matter just as long as they are on opposite sides of the equals sign

hero (hero):

Yeah, see that's what I mean. You did good by stating a "wholeness" or a "goal" to achieve, but you kinda skipped explaining "distributive rule".

OpenStudy (anonymous):

then you just go from there and move everything over then you can simplify what x is but when your moving things to the other side, make sure you change the sign so for example if its +6 and you need to move it to the other side make it negative and vise versa

hero (hero):

I can imagine that the student would have so many questions after such a short explanation

OpenStudy (anonymous):

2(x + 6) - 10 = 12 + 4(x - 1) You always do parentheses first by using the distributive property... leaving you with 2x + 12 - 10 = 12 + 4x - 4 Add like terms... x's with x's.. numbers with numbers.. I like to put the x's first and then the numbers... leaving you with 2x + 2 = 4x + 8 I like to get rid of the smallest x by taking the opposite of the smallest x.. since 2x is smaller than 4x we would get rid of the 2x by -2x from both sides. ... leaving you with 2 = 2x + 8 Then you always go to the side with the x and get rid of the number being added or subtracted from the x by taking its opposite.. so in this case you would -8 from both sides. .. leaving you with -6 = 2x Go to the side with the x and either multiply or divide... in this case divide both sides by 2... leaving you with -3 = x

hero (hero):

You're assuming the person knows what distributive property and like terms are